OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 4x + 3y = 12?

OpenStudy (across):

4x+3y=12 3y=12-4x y=4-(4/3)x The slope is -4/3

OpenStudy (karatechopper):

get y alone slope is -4/3